a child pushes a toy cart from rest on a smooth horizontal surface with a force F=5N, directed at an angle ø =10° below the horizontal. calculate the work done by the child in 5s if the carts mass is m=5kg
1
Expert's answer
2019-07-12T09:08:41-0400


Horizontal force = "F\ cos\phi=5 \ cos 10\degree = 4.92\ N"

If "a" be the horizontal acceleration then

"ma=horizontal \ force\\
5a=4.92\\
a=0.984\ m/s^2"

Distance traveled in t seconds with uniform acceleration "a" is given by

"s=ut+\frac{1}{2}at^2"

Here u= initial velocity = 0 (since the cart start from rest)

Putting t = 5 s we get

"s=0+\frac{1}{2}\times 0.984\times 5^2 \ m\\
s=12.3\ m"

Therefore, work done in 5 s is W= Horizontal force x Horizontal displacement

W= 4.92 x 12.3 J

W = 60.5 J

Answer: Work done in 5 s is 60.5 J